# ParCo: Part-Coordinating Text-to-Motion Synthesis

### v1

@@ Abstract
We study a challenging task: text-to-motion synthesis, aiming to generate motions that align with textual descriptions and exhibit coordinated movements. Currently, the part-based methods introduce part partition into the motion synthesis process to achieve finer-grained generation. However, these methods encounter challenges such as the lack of coordination between different part motions and difficulties for networks to understand part concepts. Moreover, introducing finer-grained part concepts poses computational complexity challenges. In this paper, we propose Part-Coordinating Text-to-Motion Synthesis (ParCo), endowed with enhanced capabilities for understanding part motions and communication among different part motion generators, ensuring a coordinated and fined-grained motion synthesis. Specifically, we discretize whole-body motion into multiple part motions to establish the prior concept of different parts. Afterward, we employ multiple lightweight generators designed to synthesize different part motions and coordinate them through our part coordination module. Our approach demonstrates superior performance on common benchmarks with economic computations, including HumanML3D and KIT-ML, providing substantial evidence of its effectiveness. Code is available at https://github.com/qrzou/ParCo.
@@ 4.2 Comparisons to State-of-the-art
We compare our ParCo with other methods (including the methods using ground-truth motion length) on HumanML3D test set (Table. LABEL:tab:humanml3d) and KIT-ML test set (Table. LABEL:tab:kit). Our method demonstrates superior performance compared to previous state-of-the-art methods on R-Precision and MM-Dist, and comparable results on FID, indicating ParCo’s superiority. As for Top-1, Top-2, and Top-3 of R-Precision, our ParCo surpasses previous SOTA, ReMoDiffuse [68] (using GT motion length), with 0.005, 0.008, and 0.006 on HumanML3D, 0.003, 0.008, and 0.007 on KIT-ML. And for MM-Dist, our ParCo exceeds ReMoDiffuse’s performance with 0.047 on HumanML3D and has comparable result on KIT-ML. In terms of FID, our result is on par with ReMoDiffuse on HumanML3D, and achieve the SOTA performance on HumanML3D and KIT-ML compared to the methods not relying GT information. Furthermore, our ParCo yields a lower MModality value compared to the previous state-of-the-art. This may be attributed to our ParCo generating motions that align more accurately with the text, thereby reducing the occurrence of irrelevant or incorrect motions. Qualitative results presented in Fig. 5 demonstrate that our generated motions are more realistic, coordinated, and aligned with textual descriptions for texts involving multiple body parts. Details such as "steps back," "tilts far to the left," and "stumbles" are accurately captured by our ParCo, while other methods either ignore or incorrectly generate these nuanced actions. Besides, as illustrated in Fig. LABEL:tab:humanml3d and Fig. LABEL:tab:humanml3d_trainvaltest_analysis, ParCo is endowed the part-level composition superiority across action-level spatial and temporal compositions respectively More details are provided in the supplementary materials.
@@ 4.3 Analysis
As illustrated in Table. LABEL:tab:humanml3d and Table. LABEL:tab:kit, our performance surpasses that of real motion. We attribute this phenomenon to the widely used evaluation protocol [14]. This evaluation protocol relies on pretrained feature extractors to extract meaningful features of text and motion from their raw input. However, these feature extractors are trained on the train split, lacking the knowledge of the test split data distribution. Therefore, we posit that for motions similar to those in the train split, it is easier for the extractor to extract the essential features. As demonstrated in Table. LABEL:tab:humanml3d_trainvaltest_analysis, we conduct evaluations for the real motions from Train, Val, and Test sets, respectively. The results demonstrate that the current evaluation protocol yields significantly higher Top-1/2/3 and MM-Dist metrics on the Train split compared to Val and test splits. This suggests that the pretrained feature extractors understand the data in the Train split better than the Val and Test splits. Given that our motion generation model is trained on the train split, its generated motions from any input text resemble those real data in the train split rather than the test split. And our ParCo is capable of synthesizing realistic and coordinated motions. These enables the feature extractors to capture essential information easier, making the matching of synthetic motions easier than matching of the real motions from test set. Consequently, our performance on the test set exceeds that of real motion. In addition, given the current lack of a comprehensive metric that can holistically reflect the degree of text-to-motion semantic alignment, the fidelity of motion generation, and diversity, we advocate for increased attention within the community for a more holistic assessment of text-to-motion synthesis.

### v2

@@ Abstract
We study a challenging task: text-to-motion synthesis, aiming to generate motions that align with textual descriptions and exhibit coordinated movements. Currently, the part-based methods introduce part partition into the motion synthesis process to achieve finer-grained generation. However, these methods encounter challenges such as the lack of coordination between different part motions and difficulties for networks to understand part concepts. Moreover, introducing finer-grained part concepts poses computational complexity challenges. In this paper, we propose Part-Coordinating Text-to-Motion Synthesis (ParCo), endowed with enhanced capabilities for understanding part motions and communication among different part motion generators, ensuring a coordinated and fined-grained motion synthesis. Specifically, we discretize whole-body motion into multiple part motions to establish the prior concept of different parts. Afterward, we employ multiple lightweight generators designed to synthesize different part motions and coordinate them through our part coordination module. Our approach demonstrates superior performance on common benchmarks with economic computations, including HumanML3D and KIT-ML, providing substantial evidence of its effectiveness. Code is available at: https://github.com/qrzou/ParCo.
@@ 4.2 Comparisons to State-of-the-art
We compare our ParCo with other methods (including the methods using ground-truth motion length) on HumanML3D test set (Table. LABEL:tab:humanml3d) and KIT-ML test set (Table. LABEL:tab:kit). Our method demonstrates superior performance compared to previous state-of-the-art methods on R-Precision and MM-Dist, and comparable results on FID, indicating ParCo’s superiority. As for Top-1, Top-2, and Top-3 of R-Precision, our ParCo surpasses previous SOTA, ReMoDiffuse [69] (using GT motion length), with 0.005, 0.008, and 0.006 on HumanML3D, 0.003, 0.008, and 0.007 on KIT-ML. And for MM-Dist, our ParCo exceeds ReMoDiffuse’s performance with 0.047 on HumanML3D and has comparable result on KIT-ML. In terms of FID, our result is on par with ReMoDiffuse on HumanML3D, and achieve the SOTA performance on HumanML3D and KIT-ML compared to the methods not relying GT information. Furthermore, our ParCo yields a lower MModality value compared to the previous state-of-the-art. This may be attributed to our ParCo generating motions that align more accurately with the text, thereby reducing the occurrence of irrelevant or incorrect motions. Qualitative results presented in Fig. 5 demonstrate that our generated motions are more realistic, coordinated, and aligned with textual descriptions for texts involving multiple body parts. Details such as "steps back," "tilts far to the left," and "stumbles" are accurately captured by our ParCo, while other methods either ignore or incorrectly generate these nuanced actions. Besides, as illustrated in Fig. LABEL:tab:humanml3d and Fig. LABEL:tab:humanml3d_trainvaltest_analysis, ParCo is endowed the part-level composition superiority across action-level spatial and temporal compositions respectively. More details are provided in the supplementary materials.
@@ 4.3 Analysis
As illustrated in Table LABEL:tab:humanml3d and Table LABEL:tab:kit, our performance surpasses that of real motion due to the evaluation protocol [14], which uses pretrained feature extractors trained only on the train split. These extractors are more effective at extracting features from motions similar to those in the train split. Table LABEL:tab:humanml3d_trainvaltest_analysis shows that evaluations on real motions from the Train, Val, and Test sets reveal significantly higher Top-1/2/3 and MM-Dist metrics on the Train split. This indicates better extractor understanding of the Train split data. Our model, trained on the train split, generates motions resembling the train data, making feature extraction and matching easier, leading to superior performance on the test set. Due to the lack of a comprehensive metric for text-to-motion semantic alignment, motion generation fidelity, and diversity, we advocate for a more holistic evaluation within the community.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Key R&D Program of China under Grant 2018AAA0102801.
Appendix 0.A Whole-body to Part Motions Discretization
The HumanML3D [14] and KIT-ML [48] datasets utilize the SMPL [35] and MMM [60] Human Models, respectively. These datasets include joints related to whole-body motion, excluding hand joints, as depicted in Fig. 9 and Fig. 9. In addition, HumanML3D utilizes 22 joints from the SMPL human model, while the widely-used preprocessed KIT-ML benchmark, provided by [14], comprises 21 joints.
0.A.0.1 ParCo’s 6-Part Division
Our ParCo divides the whole body into six parts: R.Leg, L.Leg, R.Arm, L.Arm, Backbone, and Root. Specific partitioning details for HumanML3D and KIT-ML are illustrated in Fig. 9 and Fig. 9. Both R.Arm and L.Arm include the 9-th joint for HumanML3D (3-th joint for KIT-ML). The inclusion of the joint in both arms is due to its role as a key point connecting the arms and the backbone, providing positional information for the arms relative to this connection point. When reconstructing the whole-body motion from part motions, we obtain three predictions of this joint from R.Arm, L.Arm, and Backbone. We use the average of these three values as the final prediction.
0.A.0.2 Upper and Lower Body Division
The upper-and-lower-body division is proposed by SCA [13], which divides the human body into upper and lower halves, both containing the backbone joints. In our ablation experiments, we perform the upper-and-lower-body division on the HumanML3D as,
  • Upper: 9, 14, 17, 19, 21, 13, 16, 18, 20, 0, 3, 6, 12, 15
  • Lower: 0, 2, 5, 8, 11, 1, 4, 7, 10, 3, 6, 9, 12, 15
where the numbers denote the joint number. It is noteworthy that, despite SCA dividing whole-body into upper and lower body for motion generation, its generations of upper body motion and lower body motion are entirely independent and lack coordination compared to our method.
Figure 8: ParCo’s 6-Part Division for SMPL Human Model.
Figure 9: ParCo’s 6-Part Division for MMM Human Model.
Appendix 0.B Details of Text-Length-Based Splits
In order to investigate the synthetic performance given textual descriptions of different lengths, we divide the HumanML3D test set into four splits based on the length of textual descriptions. The test set contains a total of 4,384 motions, each motion is described by multiple textual descriptions. Following  [67] and  [14], we set the maximum motion length to 196 and the minimum length to 40, resulting in a total of 12,635 motion-text pairs. The distribution of these pairs, sorted by text length, is shown in Fig. 0.B. We further divide these pairs into four subsets (0-25%, 25-50%, 50-75%, 75-100%) from short to long. The details are shown in Table. 0.B, including the shortest/longest text lengths, the average length, the number of pairs, and the percentage.
Table 6: VQ-VAE Reconstruction Performance on HumanML3D and KIT-ML test sets.
Datasets Methods R-Precision $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\rightarrow$
Top-1 Top-2 Top-3
HumanML3D Real Motion $0.511^{\pm.003}$ $0.703^{\pm.003}$ $0.797^{\pm.002}$ $0.002^{\pm.000}$ $2.974^{\pm.008}$ $9.503^{\pm.065}$
T2M-GPT ${0.501^{\pm.002}}$ ${0.692^{\pm.002}}$ ${0.785^{\pm.002}}$ ${0.070^{\pm.001}}$ ${{3.072^{\pm.009}}}$ ${9.593^{\pm.079}}$
Up&Low ${0.488^{\pm.002}}$ ${0.683^{\pm.002}}$ ${0.780^{\pm.002}}$ ${0.066^{\pm.001}}$ ${{3.100^{\pm.007}}}$ $\bm{9.581^{\pm.062}}$
ParCo (Ours) $\bm{0.503^{\pm.003}}$ $\bm{0.693^{\pm.003}}$ $\bm{0.790^{\pm.002}}$ $\bm{0.021^{\pm.000}}$ $\bm{3.019^{\pm.007}}$ ${9.411^{\pm.086}}$
KIT-ML Real Motion $0.424^{\pm.005}$ $0.649^{\pm.006}$ $0.779^{\pm.006}$ $0.031^{\pm.004}$ $2.788^{\pm.012}$ $11.08^{\pm.097}$
T2M-GPT $0.399^{\pm.005}$ $0.614^{\pm.005}$ $0.740^{\pm.006}$ $0.472^{\pm.011}$ $2.986^{\pm.027}$ $\bm{10.994^{\pm.120}}$
ParCo (Ours) $\bm{0.407^{\pm.007}}$ $\bm{0.629^{\pm.005}}$ $\bm{0.760^{\pm.004}}$ $\bm{0.311^{\pm.006}}$ ${\bm{2.892^{\pm.016}}}$ ${10.987^{\pm.081}}$
Table 7: Statistics of Text-Length-Based Splits.
Statistics $0-25\%$ $25-50\%$ $50-75\%$ $75-100\%$
Min length $4$ $8$ $11$ $16$
Max length $7$ $10$ $15$ $72$
Avg length $6.0$ $8.9$ $12.8$ $22.3$
Total Count $3210$ $2936$ $3096$ $3294$
Percentage(%) $25.6$ $23.4$ $24.7$ $26.3$
Figure 10: Distribution of Counts for Text Length.
Appendix 0.C VQ-VAE Reconstruction Performance
The reconstruction performance of VQ-VAEs is presented in Table. 6. Specifically, we integrate the reconstructions of part motions into the whole-body motion for evaluation. We conduct the ablation study of reconstruction performance with different partitioning methods on HumanML3D. The results indicate that the performance of our ParCo’s 6 small VQ-VAE for part motion reconstruction surpasses the upper-and-lower-body division, and outperforms the baseline [67] which employs a large-parameter VQ-VAE for whole-body motion.
Appendix 0.D Additional Training Details
During training VQ-VAE, we employ the velocity reconstruction auxiliary loss to assist training, following T2M-GPT [67] and SCA [13]. We use the last training checkpoint of VQ-VAE for the subsequent transformer’s training. For the transformer, we select the checkpoint with the lowest FID during training for text-to-motion evaluation. Additionally, we use the decoder of transformer [63] as our text-to-motion generator. The decoder of transformer achieves autoregressive prediction by masking the upper triangle of the self-attention map. To enhance the robustness of synthesis, we utilize the Corrupted Sequence [67] strategy to augment motion sequences. Inspired by MAE [19], we also introduce masked part modeling, a conceptually simple yet effective approach, to enhance part relation learning for coordinated motion generation. Specifically, we randomly replace a portion of body parts at each moment with mask tokens and force the remaining parts to predict them. Our ParCo is trained on a single A100 GPU for a total duration of 72.8 hours (20.5 hours for stage 1 and 52.3 hours for stage 2).
Figure 11: Additional qualitative comparison with existing methods. Green indicates the motion is consistent with the text description. Red indicates the text description lacks the corresponding motion or got the wrong motion.
Appendix 0.E Additional Qualitative Results
Additional qualitative results are presented in Fig. 11. The motions are generated according to text prompts from HumanML3D test set. These results demonstrate that our method can generate realistic and coordinated motions aligned with the text.